OpenStudy (anonymous):

Solve 3 (1/2 (x+5.80))=246.20

OpenStudy (freckles):

you can isolate the 3*1/2 is 3/2 So you have 3/2 (x+5.8)=246.2 multiply both sides by 2/3 and then subtract 5.8 on both sides

OpenStudy (anonymous):

Would it make a difference if the equation was 3 [1/2 (x+5.80)]=246.20

OpenStudy (freckles):

3*1/2 is still 3/2

OpenStudy (freckles):

this is the same equation you just had above

OpenStudy (freckles):

\[3[\frac{1}{2}(x+5.8)]=246.2 \\ (3 \cdot \frac{1}{2})(x+5.8)=246.2 \text{ since multiplication is commutative } \\ \frac{3}{2}(x+5.8)=246.2\]

OpenStudy (anonymous):

I am just trying to figure how I solve this to equal 94.5

OpenStudy (freckles):

well have you tried mulyiplying 2/3 on both sides yet?

OpenStudy (freckles):

multiplying*

OpenStudy (freckles):

and then finally subtract 5.8 on both sides

OpenStudy (freckles):

also 94.5 is not the right solution

OpenStudy (anonymous):

Then can you please tell me where I am going wrong on this question... Four employees are cleaning out wells. Tracy collects 5.80 more than Akuka. Kyle collects half as much as Tracey. Bev collects 3 times as much as Kyle. How much does each collect if the total is 246.20. I have... Asuka = x Tracy = x+5.8 Kyle 1/2 (x+5.8) Bev 3 (1/2 (x+5.80)

OpenStudy (freckles):

how much does each collect if the total is 246.2 you only considered bev's collectings

OpenStudy (freckles):

\[\text{ Asuka}+\text{ Tracy}+\text{ Kyle }+\text{ Bev}=246.2 \\ x+(x+5.8)+\frac{1}{2}(x+5.8)+\frac{3}{2}(x+5.8)=246.2 \\ \text{ I'm going \to factor out } (x+5.8) \text{ from last three terms } \\ x+1(x+5.8)+\frac{1}{2}(x+5.8)+\frac{3}{2}(x+5.8)=246.2 \\ x+(x+5.8)[1+\frac{1}{2}+\frac{3}{2}]=246.2 \\ x+(x+5.8)[3]=246.2 \\ x+3(x+5.8)=246.2\]

OpenStudy (freckles):

also solving for x will give you how much Asuka got and then find x+5.8 will be Tracy's amount and so on...

OpenStudy (freckles):

and judging by the answers I got for each after finding x you are looking for how much bev collected which is 94.5

OpenStudy (freckles):

anyways let me know if you have a problem solving the above equation for x and then inserting it into Bev's formula

OpenStudy (freckles):

do you understand?